OpenStudy (anonymous):

Find the f^-1(y) inverse of F(x) F(x)=x-10

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

so f^-1(x) =x+10

OpenStudy (anonymous):

Yup lol

OpenStudy (anonymous):

Oh hold on you're finding f^-1(y)

OpenStudy (anonymous):

i dont get it?....

OpenStudy (anonymous):

So since you're solving for f^-1(y) it's even easier, you're just solving for x, let f(x) = y, so y = x-10 and solve for x.

OpenStudy (anonymous):

so whats the answer.

OpenStudy (anonymous):

y = x-10 solve for x and then let x = f^-2(y)

OpenStudy (anonymous):

f^-1(y)*

OpenStudy (anonymous):

your getting me confused you tell me to slove for x so its gonna be x=y-10 all over again.

OpenStudy (anonymous):

F(x)=x-10 y=x-10 x=y-10 y=x-10 correct? or wrong final answer F^-1= 10 +x or why i dont know

OpenStudy (anonymous):

Your question is asking for f^-1(y) I gave you steps to solve for f^-1(x), but I told you it's even easier if you need to solve for f^-1(y) since you're given f(x). I think you don't know what inverse means.

OpenStudy (anonymous):

\[f^{-1}(y) \] this means you need a y variable in your final answer. You're given \[f(x) = x-10 \implies y = x-10\] now just solve for x. \[x = y+10 \implies f^{-1}(y) = y+10\] because it's a function of y we have to have a y variable or it wouldn't make sense.

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

So you have the function f(x) which we can say is going towards the right -> and the inverse of f(x) which is f^-1(x) means we go the opposite direction so towards the left <-, that's why we can cancel things out using the inverse.

OpenStudy (anonymous):

Yeah, it's sort of messy when you first learn it haha.

OpenStudy (anonymous):

what about F(x)=x/6 +3

OpenStudy (anonymous):

Are you solving for f^-1(x) or f^-1(y)?

OpenStudy (anonymous):

f^-1(y)

OpenStudy (anonymous):

So do the same steps as the last question, let f(x) = y and solve for x then let x = f^-1(y).

OpenStudy (anonymous):

F(x)=x/6 +3 y=x/6 +3 y-3= x/6 7-3/6=x so F^-1(y)= y-3/6

OpenStudy (anonymous):

Why are you dividing by 6?

OpenStudy (anonymous):

To get rid of the 6 on the right side you need to multiply both sides by 6 :p

OpenStudy (anonymous):

oh yeah my mistake so F^-1(y)=6(y-3)

OpenStudy (anonymous):

There we go :)

OpenStudy (anonymous):

F(x)=3x+7 F^-1(y)=y-7/3 correct?

OpenStudy (anonymous):

Yes, you're dividing by 3 for the whole expression right? f^-1(y) = (y-7)/3

OpenStudy (anonymous):